### v2

# BRcal: An R Package to Boldness-Recalibrate Probability Predictions

Abstract
When probability predictions are too cautious for decision making, boldness-recalibration enables responsible emboldening while maintaining the probability of calibration required by the user. We introduce BRcal, an R package implementing boldness-recalibration and supporting methodology as recently proposed. The BRcal package provides direct control of the calibration-boldness tradeoff and visualizes how different calibration levels change individual predictions. We describe the implementation details in BRcal related to non-linear optimization of boldness with a non-linear inequality constraint on calibration. Package functionality is demonstrated via a real world case study involving housing foreclosure predictions. The BRcal package is available on the Comprehensive R Archive Network (CRAN) (https://cran.r-project.org/web/packages/BRcal/index.html) and on Github (https://github.com/apguthrie/BRcal).
keywords:
Boldness, Calibration, Optimization, Probability Predictions, Forecasting, R Package
Acknowledgements The authors are thankful to Leidos for providing funding for this work. The authors would like to thank Matthew Keefe and Andrew McCoy for their insights and roles obtaining case study data for this line of research.
1 Introduction
Boldness-recalibration is a technique that responsibly emboldens probability predictions of binary events while maintaining a user chosen posterior probability of calibration [1]. When events occur at the rates that were predicted, the probability predictions are well calibrated. Calibration is important so decision makers can understand the long run rates of events. Overly cautious predictions tend to satisfy the literal definition of calibration (e.g. predicting sample proportion $\bar{y}$ for all events). When predictions exhibit greater spread, or “boldness”, they are more actionable. Boldness-recalibration leverages this tradeoff between calibration and boldness by maximizing boldness at a specified level of calibration, subject to classification accuracy. Here, calibration is measured using the posterior model probability of calibration. The posterior model probability of calibration is easily interpretable in the context of an individual forecaster, which is an advantage for a calibration metric. The key feature of boldness-recalibration that sets it apart from most other calibration techniques is that it provides responsible emboldening of predictions rather than just improving calibration. In this paper, we present BRcal: an R package implementing boldness-recalibration and supporting methodology including calibration assessment via Bayesian model selection, LLO-adjustment via MLEs, and visualizations.
Many R packages exist that implement recalibration techniques, calibration metrics, and visualizations such as reliability diagrams. The CalibrationCurves [2] and pmcalibration [3] packages both implement and visualize Cox linear logistic calibration [4]. An implementation of isotonic regression for calibration [5] can be found in the rfUtilities package [6]. The platt package [7] focuses on Platt Scaling for calibration [8]. The betacal [9] package provides Beta calibration with options for both 2- and 3-parameter approaches [10]. A bootstrap-based recalibration approach for a variety of regression models, including visualizations and scoring, can be found in the rms package. The braggR package implements the approach for improving calibration via aggregation proposed in [11]. The scoring [12], scoringutils [13], DescTools [14], and Yardstick [15] packages primarily focus on scoring rules and other metrics including the Brier Score [16], Expected Calibration Error (ECE), and Log Loss. Packages like calibration [17], Calibtools [18], CalibratR [19], and probably [20] provide a suite of calibration tools including histogram binning [21], Bayesian Binning into Quantiles [22] and several of the previously mention recalibration techniques, scoring metrics, and visualizations. None of these include boldness-recalibration.
The rest of this paper is organized as follows. Section 2 briefly reviews the methodology behind boldness-recalibration and related techniques. Section 3 introduces the BRcal package by describing its core functions and implementation. In Section 4, we demonstrate the capabilities of BRcal by presenting an example related to housing foreclosure predictions. Section 5 provides concluding remarks.
2 Methods
We introduced boldness-recalibration, a forecaster agnostic approach, to improve upon established recalibration techniques in situations where overly cautious predictions hinder the decision making process [1]. However, it is important to assess calibration and boldness before boldness-recalibrating. Rather than rely on calibration assessment techniques such as the Brier Score, Expected Calibration Error (ECE), and Log Loss, we employ a Bayesian model selection approach to measure the probability of calibration. Throughout this paper and in the BRcal package, we focus on probability predictions of binary events.
2.1 Background
For a more complete exposition of methods, see [1]. A brief and self-contained review is found here. Let $\mathbf{x}$ be a vector of probability predictions from a single source (i.e. a single forecaster, or a single machine learning model, etc) and let $\mathbf{y}$ be a vector of binary event outcomes. Following the convention of [1], we adopt the following likelihood function
$\pi(\mathbf{x},\mathbf{y}|\delta,\gamma)=\prod_{i=1}^{n}c(x_{i};\delta,\gamma)^{y_{i}}\left[1-c(x_{i};\delta,\gamma)\right]^{1-y_{i}}.$ (1)
where $c(x_{i};\delta,\gamma)$ is the Linear Log Odds (LLO) recalibration function defined as
$c(x_{i};\delta,\gamma)=\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}},$ (2)
and $\delta>0$ and $\gamma\in\mathbb{R}$ are recalibration parameters [23, 24]. Equation (2) maps probability forecasts to the log odds scale, shifts them by log $\delta$, scales by $\gamma$, and remaps back to the probability scale. When we apply (2) to prediction $x_{i}$, we call this “LLO-adjustment”.
To assess calibration, we take an approximate Bayesian model selection-based approach comparing a well calibrated model ($\delta$ = $\gamma$ = 1) denoted by $M_{c}$, to an uncalibrated model ($\delta>0$ and $\gamma\in\mathbb{R}$) denoted by $M_{u}$. The posterior model probability of $M_{c}$ is defined as
$P(M_{c}|\mathbf{y})=\frac{1}{1+BF\frac{P(M_{u})}{P(M_{c})}}.$ (3)
where $BF$ is the Bayes Factor comparing $M_{u}$ to $M_{c}$ defined as
$BF=\frac{P(\mathbf{y}|M_{u})}{P(\mathbf{y}|M_{c})}.$ (4)
We approximate the Bayes Factor using the Bayesian Information Criteria (BIC) described in [25] and [26] such that
$BF\approx exp\left\{-\frac{1}{2}(BIC_{u}-BIC_{c})\right\}.$ (5)
Here, the BIC under the well calibrated model $M_{c}$ is defined as
$BIC_{c}$ $=-2\times log(\pi(\delta=1,\gamma=1|\mathbf{x},\mathbf{y})).$ (6)
The penalty term for number of estimated parameters is omitted in (6) as both parameters are fixed at 1 under $M_{c}$. The BIC under the uncalibrated model $M_{u}$ is defined as
$BIC_{u}$ $=2\times log(n)-2\times log(\pi(\hat{\delta}_{MLE},\hat{\gamma}_{MLE}|\mathbf{x},\mathbf{y})).$ (7)
The quantity $P(M_{c}|\mathbf{y})$ is the probability $\mathbf{x}$ is calibrated given observed outcomes $\mathbf{y}$ and serves as our measure of calibration.
To recalibrate $\mathbf{x}$ such that calibration is maximized, we use (1) to obtain maximum likelihood estimates (MLEs) for the recalibration parameters, denoted $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$. Then, these MLEs are used to obtain the maximally calibrated predictions $\mathbf{x}_{MLE}$ where $x_{i,MLE}=c(x_{i};\hat{\delta}_{MLE},\hat{\gamma}_{MLE})$. We call this approach MLE-recalibration.
2.2 Boldness-Recalibration
Maximally calibrated predictions are not necessarily bold enough for decision making. For example, the Hockey study in [1] showed that even after MLE-recalibration, the expert forecaster’s predictions could be more bold while maintaining a high probability of calibration via boldness-recalibration. On the other hand, random noise predictions were reigned in close to the base rate of 0.53 (i.e. not bold) to be maximally calibrated. In general, predicting the base rate $\bar{y}$ for all events satisfies the definition of calibration, but lacks boldness. We define Boldness as the spread in probability predictions and measured by
$s_{b}$ $=sd(\mathbf{x}),$ (8)
Boldness-recalibration finds values of $\delta$ and $\gamma$ that maximize $s_{b}$ subject to a user specified level of calibration denoted by $t$. The parameters that achieve this goal are denoted by $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ and are defined as
$(\hat{\delta}_{t},\hat{\gamma}_{t})$ $=argmax_{(\hat{\delta},\hat{\gamma})}(s_{b}:P(M_{c}|\mathbf{y},\hat{\delta},\hat{\gamma})\geq t).$ (9)
Then $\mathbf{x}_{t}$ are called the $(100*t)\%$ boldness-recalibrated predictions where $x_{i,t}=c(x_{i},\hat{\delta}_{t},\hat{\gamma}_{t})$.
The boldness-recalibration methodology formulates finding $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ as a nonlinear constrained optimization problem. This optimization is nonlinear because the objective function we optimize is nonlinear with respect to parameters $\delta$ and $\gamma$ and is formulated as
$f(\delta,\gamma)$ $=-sd(\mathbf{x}^{\prime})$ (10)
$=-\sqrt{\frac{\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})^{2}}{n-1}}$ (11)
where $x_{i}^{\prime}=c(x_{i},\delta,\gamma)$ and $\bar{x}^{\prime}$ is the mean of $\mathbf{x}^{\prime}$. This is a constrained optimization problem because we restrict the region of parameter values over which we optimize using the following inequality constraint function,
$g(\delta,\gamma)$ $=-(P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})-t)\leq 0,$ (12)
[…]
To leverage the efficient gradient-based optimization routines, we also define the Jacobian of (10) as $J_{f}(\delta,\gamma)=\left[\frac{\partial f}{\partial\delta},\frac{\partial f}{\partial\gamma}\right]$ where
$\frac{\partial f}{\partial\delta}$ $=\frac{-1}{\delta(n-1)sd(\mathbf{x}^{\prime})}\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(x_{i}^{\prime}(1-x_{i}^{\prime})-\frac{1}{n}\sum_{j=1}^{n}x_{j}^{\prime}(1-x_{j}^{\prime})\right),$ (13)
$\frac{\partial f}{\partial\gamma}$ $=\frac{-1}{(n-1)sd(\mathbf{x}^{\prime})}\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime}(1-x_{i}^{\prime})-\frac{1}{n}\sum_{j=1}^{n}log\left(\frac{x_{j}}{1-x_{j}}\right)x_{j}^{\prime}(1-x_{j}^{\prime})\right),$ (14)
and the Jacobian of (12) as $J_{g}(\delta,\gamma)=\left[\frac{\partial g}{\partial\delta},\frac{\partial g}{\partial\gamma}\right]$ where
$\frac{\partial g}{\partial\delta}$ $=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\frac{1}{\delta}\left(\sum_{i=1}^{n}\hat{\gamma}_{MLE}^{\prime}(y_{i}-x_{i,MLE}^{\prime})-y_{i}+x_{i}^{\prime}\right),$ (15)
$\frac{\partial g}{\partial\gamma}$ $=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)\big{(}\hat{\gamma}_{MLE}^{\prime}(y_{i}-x_{i,MLE}^{\prime})-y_{i}+x_{i}^{\prime}\big{)}\bigg{)}.$ (16)
Here, $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ are the MLEs for $x_{i}^{\prime}$ and $x_{i,MLE}^{\prime}=c(x_{i}^{\prime};\hat{\delta}_{MLE}^{\prime},\hat{\gamma}_{MLE}^{\prime})$. Full derivation of these gradients can be found in Appendix A.
3 The BRcal package
3.1 Overview
The main contributions of the BRcal package are (i) boldness-recalibration, (ii) MLE-recalibration, (iii) calibration assessment, and (iv) supporting visualizations. The brcal() function facilitates boldness-recalibration at a user-specified level of calibration, The mle_recal() function facilitates MLE-recalibration. The bayes_ms() and llo_lrt() functions facilitate Bayesian and frequentist approaches to calibration assessment, respectively. The plot_params() function provides a visualization for the constrained optimization via a color contour plot of the posterior model probability in (12). The lineplot() function provides a visualization for examining how individual predictions change under boldness-recalibration and MLE-recalibration (optionally). The package also provides a supporting function, LLO(), for LLO-adjusting via any specified parameters.
The BRcal package has three main dependencies on other R packages. The brcal function requires the nloptr package for optimization [27]. Additionally, the plotting functions make use of the fields [28] and ggplot2 packages [29].
The BRcal package is available on the Comprehensive R Archive Network (CRAN) (https://cran.r-project.org/package=BRcal) and on Github (https://github.com/apguthrie/BRcal). For additional information and a helpful user guide with a tutorial, see the package vignette included on CRAN.
3.2 Data
The BRcal package includes two built-in datasets called foreclosure and hockey. The foreclosure dataset includes probability predictions of foreclosure and the foreclosure status of 5,000 randomly selected housing transactions in 2010 from Wayne County, Michigan. These data arose from a project that developed a monitoring strategy for the housing market using a spatially-adjusted model of foreclosure risk [30]. The probability of foreclosure was estimated using a mixture model of spatial kernel density estimates. The model was trained during the first six months of 2005 using data from Wayne County, Michigan, which is a pre-Great Recession time period. The foreclosure dataframe includes three columns:
  • y - integer; sale type, 1 = foreclosure, 0 = regular sale
  • x - numeric; predicted probabilities of foreclosure
  • year - numeric; year of observed foreclosure or regular sale
The hockey dataset includes probability predictions of a home team win from FiveThirtyEight [31] and the winner of each game in the 2020-2021 National Hockey League (NHL) season. These data were used in [1], and more information about the data can be found there. The hockey data frame includes four columns:
  • y - integer; game result, 1 = home team win, 0 = home team loss
  • x - numeric; predicted probabilities of a home team win
  • rand - numeric; randomly generated predicted probabilities of a home team win
  • winner - factor; game result, “home” = home team win, “away” = home team loss.
3.3 The brcal() Function
The core function of the BRcal package is the brcal() function, which implements boldness-recalibration. It takes two required arguments, x a vector of probability predictions, and y a vector of binary event outcomes. The other key argument is t, which represents the minimum desired probability of calibration as seen in (9) and is set to 0.95 by default. Users should adjust this setting based on the level of risk they are willing to accept by sacrificing some calibration in favor of increased boldness. The arguments for brcal() are summarized in Table 1.
Function Arguments Description
brcal() x Vector of probability predictions
y Vector of binary event outcomes
t=0.95 Level of user required calibration probability between 0 and 1
Pmc=0.5 Prior model probability for calibrated model $M_{c}$
tau=FALSE Calculates in terms of $\tau=log(\delta)$ instead of $\delta$ when TRUE
event=1 Value in y that represents an “event”
start_at_MLEs=TRUE Optimization starts at MLEs when TRUE
x0=NULL Starting parameter values if start_at_MLEs = FALSE
lb=c(1e-05, -Inf) Lower bounds for $\delta$ and $\gamma$
ub=c(Inf, Inf) Upper bounds for $\delta$ and $\gamma$
maxeval=500 Max number of iterations for inner optimization
maxtime=NULL Max evaluation time (in seconds)
xtol_rel_outer=1.0e-6 Tolerance for relative difference between iterations of outer optimization
xtol_rel_inner=1.0e-6 Tolerance for relative difference between iterations of inner optimization
print_level=3 Controls the amount of output from nloptr()
epsilon=2.220446e-16 Amount by which probabilities in x are pushed away from 0 or 1 from numerical stability
opts=NULL List with options to be passed to nloptr()
optim_options=NULL List with options to be passed to optim()
Table 1: Description of function arguments and their default values for the brcal() function.
The brcal() function uses the optimization capabilities of nloptr() for the nonlinear constrained optimization to find $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$ and optim() to find $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$. We primarily focus on how brcal() uses the nloptr() function in this section. Discussion of optim() to find the MLEs, which are used in boldness-recalibration, can be found in Section 3.4. The nloptr package implements both global and local optimization routines and both gradient-based and derivative-free approaches. For all sets of predicted probabilities the authors encountered or generated, the constraint surface (the posterior model probability) has been unimodal. Within the constrained region, the objective surface (the standard deviation) has also been unimodal. With this in mind, we chose to focus on local gradient based approaches of which there were three options that supported non-linear inequality constraints: Sequential Least-Squares Quadratic Programming (SLSQP) [32, 33], Augmented Lagrangian algorithm (AUGLAG) [34, 35], and Method of Moving Asymptotes (MMA) [36]. Of these, AUGLAG produced the most stable results in terms of convergence in boldness-recalibration problems and is set as the default optimization routine in brcal().
The implementation of AUGLAG provided by nloptr reformulates the constrained optimization problem as a two-stage optimization. The inner stage optimizes the objective function with a penalty for violating the constraint rather than use the constraint directly. The outer stage governs the penalty. If the inner stage returns a value that violates the constraint despite the penalty, the outer stage imposes a larger penalty. Note that since the inner stage is itself an unconstrained optimization routine, users may specify any optimization routine that supports nonlinear objective functions. For the purpose of this package, we found that SLSQP as the inner optimization routine in AUGLAG provided the most stable results, thus SLSQP is set as the default inner optimization routine.
As with any optimization problem, it is important to verify convergence and adjust arguments as needed to achieve convergence. The brcal() function provides both of these by allowing, but not requiring, full adjustment of the nloptr() and optim() routines, and returning important convergence information. The default arguments used in brcal() have been effective in most applications explored by the authors. A few options for stopping criteria, starting values, parameter bounds, and level of output are specified directly in brcal(). All other nloptr() arguments can be optionally adjusted by passing them as a list to the opts argument. Similarly, all arguments in optim() can be adjusted by passing them as a list to optim_options.
With both an inner and outer optimization routine, stopping criteria are specified for each routine separately. The nloptr package offers a variety of options including stopping based on the computation time, number of iterations, or change in function value from previous iteration. By default, the brcal() function stops the outer optimization when the relative difference in function value is $\leq 1.0e-6$ and stops the inner optimization when either the relative difference in function value is $\leq 1.0e-6$ or the number of evaluations surpasses 500. Users of brcal() are given the option to tweak any of these three options via parameters xtol_rel_outer, xtol_rel_inner, and maxevel, or users can set the max evaluation time (in seconds) using maxtime. All other stopping criteria can be set using opts.
To improve convergence, we set the default starting location of the nloptr() optimizer to the MLEs for the original probability predictions. The constraint function in (12) is maximized at the MLEs, and since the constraint surface has been empirically unimodal, these MLEs lie at the center of the constraint region. This is illustrated in Figure 1 by the $\times$ denoting the MLEs located at the center of the shaded region. By starting at the center of the plausible region of parameter values, this prevents the optimizer from being unable to locate the region of interest. We allow users to toggle this option via the start_at_MLEs parameter. By default, this parameter is set to TRUE. If set to FALSE, users can specify an alternative starting value via argument x0. Should users specify their own x0, the authors recommend imposing bounds on the parameter space using lb and ub to set lower and upper bounds. respectively, on $\delta$ and $\gamma$.
3.4 Calibration Assessment
While the goal of boldness-recalibration is to embolden when predictions are accurate, users can first assess calibration using the bayes_ms() and llo_lrt() functions. Both functions provide MLEs for $\delta$ and $\gamma$. The bayes_ms() function uses the Bayesian model selection-based approach described in Section 2.1. In addition to MLEs, this function returns the posterior model probability of calibration calculated via (3), and the BIC for both $M_{c}$ and $M_{u}$ in (6) and (7). The llo_lrt() function conducts the likelihood ratio test-based approach to assess calibration, details can be found in [1]. In addition to MLEs, this function returns the p-value associated with the test, and the test statistic. Details of the arguments for bayes_ms(), llo_lrt(), and other supporting functions in the BRcal package can be found in Table 2.
Function Arguments Description
bayes_ms() x Vector of probability predictions
llo_lrt() y Vector of binary event outcomes
Pmc=0.5 Prior model probability for calibrated model $M_{c}$ (bayes_ms() only)
event=1 Value in y that represents an “event”
optim_details=TRUE Returns optimization details from optim() when TRUE
epsilon=2.220446e-16 Amount by which probabilities in x are pushed away from 0 or 1 from numerical stability
... Additional arguments to be passed to optim()
mle_recal() x Vector of probability predictions
y Vector of binary event outcomes
probs_only=TRUE Only returns the vector of recalibrated probabilities when TRUE
event=1 Value in y that represents and “event”
optim_details=TRUE Returns optimization details from optim() when TRUE
... Additional arguments to be passed to optim()
LLO() x Vector of probability predictions
delta Value for $\delta$ to LLO-adjust x by
gamma Value for $\gamma$ to LLO-adjust x by
Table 2: Description of the function arguments and default values for the bayes_ms(), llo_lrt(), mle_recal(), and LLO() functions in the BRcal package.
Both bayes_ms() and llo_lrt() rely on a call to the optim() function to maximize the log of the likelihood in (1) to find $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$. By default optim() minimizes the function it is given. To follow this convention, we choose to minimize the negative of the log likelihood. Additionally, optim() offers several options for optimization routines. Given that parameter space of $\gamma$ is unbounded but $\delta>0$, we choose to re-parameterize the log likelihood for sake of optimization. Instead of operating on $\delta$, we use $\tau=log(\delta)\in\mathbb{R}$ [23]. With this version of the log likelihood, we choose to use the default routine, "Nelder-Mead" [37]. While by default we operate on $\tau$ instead of $\delta$ in the optimization step, bayes_ms() and llo_lrt() also have a parameter called tau which will change the output of the function to be in terms of $\tau$ instead of $\delta$. To make any adjustments to the call to optim(), users can pass any arguments to optim() using ....
3.5 MLE-Recalibration
The BRcal package also provides mle_recal() to maximally calibrate predictions by LLO-adjusting via the MLEs. This function returns $\mathbf{x}_{MLE}$, where $x_{MLE,i}=c(x_{i};\hat{\delta}_{MLE},\hat{\gamma}_{MLE})$, for a given set of probability predictions $\mathbf{x}$ and corresponding outcomes. Users do not need to specify the MLEs as the function calculates this for them.
Should users want to LLO-adjust based on parameter values other than the MLEs or $(100*t)\%$ boldness-recalibration parameters, we also provide the LLO() function. This function takes argument x as the vector of probabilities to which the LLO function will be applied. Arguments delta and gamma correspond to $\delta$ and $\gamma$, respectively, in (2). The returned vector of probabilities is $\mathbf{x}^{\prime}$ where $x^{\prime}_{i}=c(x_{i};\delta^{\prime},\gamma^{\prime}$), and $\mathbf{x},\delta^{\prime},\gamma^{\prime}$ are the user provided arguments.
3.6 Visualization Functions
The BRcal package provides two types of visualizations to conduct and assess the impact of boldness-recalibration: (i) a lineplot visualizing how individual predictions change under various LLO-adjustments and (ii) a color image and contour plot of the posterior model probability as a function of $\delta$ and $\gamma$. Examples of both are provided in Figures 2 and 3 in Section 4. This section describes the implementation details for these plot types. Description of the plotting function arguments and default values can be found in Table 3.
Function Arguments Description
lineplot() x Vector of probability predictions
y Vector of binary event outcomes
t_levels=NULL Vector of desired level(s) of calibration at which to boldness-recalibrate
plot_original=TRUE Plots the original probabilities when TRUE
plot_MLE=TRUE Plots the MLE Recalibrated probabilities when TRUE
df=NULL Dataframe of adjusted probabilities returned by previous call to lineplot()
return_df=FALSE Returns specially formatted dataframe of LLO-adjusted probabilities when TRUE
thin_to=NULL Randomly thins to the specified number of observations
thin_prop=NULL Randomly thins to the specified proportion
thin_by=NULL Thins by taking every thin_by observations
seed=0 Seed used for random thinning for repeatability
ggpoint_options=NULL List of additional arguments passed to geom_point()
ggline_options=NULL List of additional arguments passed to geom_line()
plot_params() x Vector of probability predictions
y Vector of binary event outcomes
z=NULL Matrix of posterior model probabilities across $k$ by $k$ grid of $\delta$ and $\gamma$ values returned by previous call to plot_params()
t_levels=NULL Vector of desired level(s) of calibration at which to plot contours
k=100 Grid of $\delta$ and $\gamma$ values is size $k$ by $k$
dlim=c(1e-04,5) Bounds for $\delta$
glim=c(1e-04,5) Bounds for $\gamma$
return_z=FALSE Returns matrix of posterior model probabilities when TRUE
contours_only=FALSE Only plots contours at levels in t_levels when TRUE with no grid cell coloring
imgplt_options=list(...) List of additional arguments passed to image.plot()
contour_options=list(...) List of additional arguments passed to image.plot()
Table 3: Description of primary function arguments and their default values for the visualization functions (lineplot() and plot_params()) in the BRcal package. Note not all arguments are listed here. See package documentation and vignette on CRAN for more details.
3.6.1 Lineplot
The first visualization provided is a lineplot showing how probability predictions x change after LLO-adjustment implemented in the lineplot() function and utilizes ggplot2 graphics [29]. This function allows users to pass a set of predictions x and outcomes y, and specify which predictions (original, MLE-recalibrated, tboldness-recalibrated) they are interested in plotting. By default, only the original and MLE-recalibrated probabilities are plotted. To omit either the original or MLE-recalibrated sets, users can set plot_original or plot_MLE to FALSE. To specify levels of boldness-recalibration, users can pass a vector of values for $t$ via t_levels. The resulting plot shows each specified set of probabilities side-by-side with lines connecting corresponding predictions. See Figure 2 for an example. The lines are color coded based on the corresponding outcome, where blue represents $y_{i}$ = 1 (i.e. an “event”) and red represents $y_{i}$ = 0 (i.e. a “non-event”). The goal is to visualize (i) the overall change in boldness between the sets (e.g. spread increases/decreases under $95\%$ boldness-recalibration), (ii) how individual predictions change (e.g. a prediction of 0.7 is adjusted to 0.8), and (iii) if predictions with high predicted probabilities typically result in events (e.g. point/lines for higher probabilities are typically blue, indicating an “event”).
When the number of predictions for your lineplot is very large, plotting may be slow. To ease the graphical burden, we offer three thinning options to reduce the amount of predictions that are visualized. The thin_to argument allows users to specify some $n_{thin}$ = the number of predictions they would like to reduce their sample to. In this case, lineplot() will randomly select $n_{thin}$ predictions from x and their corresponding outcomes in y. Using thin_prop allows users to select the proportion of predictions to plot, which results in a random selection of that proportion size. Lastly, specifying the thin_by argument to some number $m$ where lineplot() then selects every $m$ observation from the predictions. Note that under all thinning strategies, all underlying calculations including the posterior model probabilities, boldness-recalibration parameters, and MLEs are conducted using the full set of probability predictions. Thinning only reduces the number of observations seen on the plot.
Once users have settled on the desired LLO-adjustments to be plotted (i.e. desired levels of boldness-recalibration, MLE-recalibration, etc), you can specify return_df=TRUE to return the underlying dataframe of adjusted probabilities to avoid recalculation of these values. This dataframe can then be reused by passing it to argument df, indicating to lineplot() that no recalculations are needed. For example, it is useful to reuse df so that users do not need to wait for the underlying boldness-recalibration calculations each time they make minor cosmetic adjustments to the plot, like modifying axes labels or font sizes.
Users can make additional plotting adjustments to the underlying calls to geom_point() and geom_line() by passing the appropriate arguments as a list to ggpoint_options and ggline_options(), respectively. Examples showing how to use these arguments can be found in the package vignette.
3.6.2 Color Contour Plot
The second visualization type is a color contour plot showing how the posterior model probability, $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$, changes as a function of $\delta$ and $\gamma$. This visualization is implemented in the plot_params function and utilizes base R graphics and the fields package [28]. The resulting plot shows the posterior model probability surface over a $k$ by $k$ grid of $\delta$ and $\gamma$ values using a color scale to differentiate high and low values of $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$. Additionally, users have the option to add contours at specified levels of calibration using t_levels. The goal of this is to visualize different values of $t$ at which a user may want to boldness-recalibrate, but does not perform boldness-recalibration nor identify $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$.
Some care is required for the selection of the range of $\delta$ and $\gamma$ via arguments dlim and glim, and the size of $k$ via argument k, which together determine the fineness of your grid of parameter values. Small values of k or misspecified parameter bounds may cause difficulties in detecting the region of non-zero $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$s. Large values of k may cause a computational bottleneck as computation time increases with k. Appendix B presents recommendations for specifying k, dlim, and glim. Once users have settled on values of these arguments, you can specify return_z=TRUE to return the underlying matrix of posterior model probabilities to avoid recalculation of these values. This matrix can then be reused by passing it to argument z, indicating to plot_params() that no recalculations are needed.
Users can make additional plotting adjustments to the underlying calls to image.plot() and countour() by passing the appropriate arguments as a list to imgplt_options and contour_options, respectively.
4 Application
In this section we demonstrate BRcal using a case study related to the foreclosure risk modelling presented by Keefe et al. [30]. Note that in all places in the output where ... appears, this indicates that the output was truncated for readability.
4.1 Foreclosure Monitoring Predictions Case Study
This example uses the foreclosure dataframe built into the BRcal package. We begin by loading the BRcal package and foreclosure data and printing the first few rows of the data via the code below.
> library(BRcal)
> data(foreclosure)
> head(foreclosure)
  y         x year
1 1 0.1155925 2010
2 1 0.3856198 2010
3 0 0.1273048 2010
4 1 0.4212137 2010
5 1 0.2623377 2010
6 1 0.5727253 2010
4.1.1 Assessing Calibration
With the data loaded, we can assess the calibration of the foreclosure predictions. First we will use the bayes_ms function to use the approximate Bayesian model selection approach.
> bt <- bayes_ms(foreclosure\$x, foreclosure\$y)
> bt
\$Pmc
[1] 0.5
\$BIC_Mc
[1] 8344.346
\$BIC_Mu
[1] 4159.783
\$BF
[1] Inf
\$posterior_model_prob
[1] 0
\$MLEs
[1] 11.109412  1.270519
\$optim_details
\$optim_details\$par
[1] 11.109412  1.270519
\$optim_details\$value
[1] 2071.375
\$optim_details\$counts
function gradient
      85       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
Notice from the output that bayes_ms() returns seven quantities in the form of a list. The prior model probability for $M_{c}$ is returned in Pmc to remind users what value was specified (0.5 by default). List items BIC_Mc and BIC_Mu in (6) and (7) are the BICs for the calibrated model and uncalibrated model, respectively. The Bayes factor comparing $M_{u}$ to $M_{c}$ in (5) is returned in BF and posterior_model_prob is the posterior model probability of calibration from (3). The MLEs for the vector x are returned in MLEs. Additionally, the list returned by the call to optim() to get the MLEs is returned in sublist optim_details.
We now show the likelihood ratio test for calibration implemented in llo_lrt(). This function returns the corresponding test statistic in test_stat, p-value in pval, MLEs in MLEs, and the corresponding list from optim() in optim_details. As expected, the MLEs and optimization details from llo_lrt() match those from bayes_ms().
> llo_lrt(foreclosure\$x, foreclosure\$y)
\$test_stat
[1] 4201.597
\$pval
[1] 0
\$MLEs
[1] 11.109412  1.270519
\$optim_details
\$optim_details\$par
[1] 11.109412  1.270519
\$optim_details\$value
[1] 2071.375
\$optim_details\$counts
function gradient
      85       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
The results of bayes_ms() and llo_lrt() both indicate that the foreclosure predictions from the model trained in 2005 are poorly calibrated, as the posterior model probability and the p-value are near 0.
4.1.2 MLE Recalibration
Next, we demonstrate how to use mle_recal() to recalibrate the foreclosure predictions such that calibration is maximized. By default, mle_recal() returns a list with (i) the vector of maximally calibrated predictions in probs, (ii) the MLEs in MLEs and (iii) the details from optim() in optim_details. Should a user only want the vector of MLE-recalibrated probabilities, they can set the option probs_only = TRUE.
> mle <- mle_recal(foreclosure\$x, foreclosure\$y, probs_only = FALSE)
> mle
\$probs
[1] 0.4557362 0.8600894 0.4905113 0.8812196 0.7491834 0.9415870
[7] 0.9247097 0.5139758 0.9206578 0.6645239 0.5170544 0.8372189
[997] 0.8972962 0.8860154 0.3612379 0.9433734
[ reached getOption("max.print") -- omitted 4000 entries ]
\$MLEs
[1] 11.109412  1.270519
\$optim_details
\$optim_details\$par
[1] 11.109412  1.270519
\$optim_details\$value
[1] 2071.375
\$optim_details\$counts
function gradient
      85       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
4.1.3 Boldness-Recalibration
Now we boldness-recalibrate using the brcal() function. By default, brcal() performs 95% boldness-recalibration and prints each step of the optimization process from nloptr(). To suppress all output from nloptr(), users can set print_level=0. Other options to minimize output can be found in the nloptr() documentation and can be passed to brcal() via print_level. Below we show a example where the full output from the first and last steps are show, the in between stages are truncated for readability.
The brcal() function returns a list with the following entries: (i) the information returned by the nloptr() call in nloptr, (ii) the prior model probability for $M_{c}$ in Pmc, (iii) the user specified level of calibration desired in t, (iv) the $(100*t)\%$ boldness-recalibration parameters obtained from nloptr() in BR_params, (v) the maximum boldness achievable under $(100*t)\%$ boldness-recalibration in sb, and (vi) the resulting probabilities from recalibrating using the $(100*t)\%$ boldness-recalibration parameters in probs.
> br95 <- brcal(hockey\$x, hockey\$y)
iteration: 1
Ψx = (11.109412, 1.270519)
Ψf(x) = -0.179135
Ψg(x) = -0.049800
iteration: 201
Ψx = (12.134803, 1.411657)
Ψf(x) = -0.199055
Ψg(x) = -0.000000
> br95
\$nloptr
Call:
nloptr::nloptr(x0 = c(11.109412, 1.270519), eval_f = obj_f, eval_grad_f =
    obj_grad_f, eval_g_ineq = constr_g, eval_jac_g_ineq = constr_grad_g,
    opts = list(algorithm = "NLOPT_LD_AUGLAG", maxeval = 500,
        maxtime = -1, xtol_rel = 1e-06, print_level = 3, local_opts =
        list(algorithm = "NLOPT_LD_SLSQP", eval_grad_f = obj_grad_f,
        eval_jac_g_ineq = constr_grad_g, xtol_rel = 1e-06)),
    t = 0.95, tau = FALSE, Pmc = 0.5, epsilon = 2.22044604925031e-16)
Minimization using NLopt version 2.7.1
NLopt solver status: 4 ( NLOPT_XTOL_REACHED: Optimization stopped
because xtol_rel or xtol_abs (above) was reached. )
Number of Iterations....: 201
Termination conditions:  maxeval: 500Ψxtol_rel: 1e-06
Number of inequality constraints:  1
Number of equality constraints:    0
Optimal value of objective function:  -0.199055078798425
Optimal value of controls: 12.1348 1.411657
\$Pmc
[1] 0.5
\$t
[1] 0.95
\$BR_params
[1] 12.134803  1.411657
\$sb
[1] 0.1990551
\$probs
   [1] 0.4069886 0.8627809 0.4448838 0.8856910 0.7381995 0.9483221
   [7] 0.9313465 0.4707687 0.9271990 0.6411276 0.4741822 0.8375914
 [997] 0.9028353 0.8908336 0.3073575 0.9500874
 [ reached getOption("max.print") -- omitted 4000 entries ]
Notice the large amount of output from nloptr() which is returned by brcal() in nloptr. As with any optimization routine, we include this copious output to allow users to check that the routine converged properly. While this output shows the value achieved optimal values of the parameters and objective function, we choose to return these separately in the list items BR_params and sb, respectively for easier extraction. The resulting set of boldness-recalibrated probabilities are also returned for easier extraction.
We enforce different restrictions on calibration by changing the value of parameter t. The two lines of code below perform 90% and 80% boldness-recalibration. We also demonstrate suppressing the output from nloptr by setting print_level=0.
br90 <- brcal(foreclosure\$x, foreclosure\$y, t=0.9, print_level=0)
br80 <- brcal(foreclosure\$x, foreclosure\$y, t=0.8, print_level=0)
Note that by leveraging the gradient information and efficient nloptr package as described in Section 3.3, each of these calls to brcal() for the foreclosure data takes approximately 60 seconds or less to run.
4.1.4 LLO-Adjustment
The LLO() function allows users to directly LLO-adjust a vector of probability predictions, passed via x, by shift parameter delta and scale parameter gamma. While most users of this package may not use the LLO() function directly, there are some cases where users may want to do so. One example of this could be when you have MLEs for $\delta$ and $\gamma$ based on historical data and want to LLO-adjust unlabeled out of sample predictions. A user could use the LLO function to LLO-adjust the new set of predictions by passing them to function argument x, and the MLEs for $\delta$ and $\gamma$. Another use case could be when you have boldness-recalibration parameter estimates from brcal() but did not save the boldness-recalibrated probabilities. Users could LLO-adjust their probabilities using the boldness-recalibration parameter estimates using LLO()without needing to re-compute them using brcal(), thus avoiding unnessesary computation time. This is demonstrated below with the 95% boldness-recalibration estimates saved in section 4.1.3.
> LLO(foreclosure\$x, delta=br95\$BR_params[1], gamma=br95\$BR_params[2])
   [1] 0.4069886 0.8627809 0.4448838 0.8856910 0.7381995 0.9483221
   [7] 0.9313465 0.4707687 0.9271990 0.6411276 0.4741822 0.8375914
 [997] 0.9028353 0.8908336 0.3073575 0.9500874
 [ reached getOption("max.print") -- omitted 4000 entries ]
Notice in the output above that we get the same boldness-recalibrated probabilities as we did using brcal() in section 4.1.3.
4.1.5 Visualizations
First we will use lineplot to visualize how MLE-recalibration, t=95, 90, and 80% boldness-recalibration changed the individual probabilities themselves. To do so, we pass the vector of the original probability predictions and the corresponding outcomes along with the levels of boldness-recalibration we want to visualize using t_levels. The code below generates the lineplot that is shown on the left in Figure 2. Given the size of this data, the plot may take a few minutes to generate. In practice, users can leverage the strategies discussed in section 3.6.1 for reducing computation time.
> lp <- lineplot(foreclosure\$x, foreclosure\$y, t_levels=c(0.95, 0.9, 0.8),
                 return_df=TRUE)
> lp\$plot
While it can be useful to visualize all of these sets at the same time, users may want to focus on how emboldening affects predictions after they have already been MLE-recalibrated. The following line of code uses the dataframe returned from the first call to lineplot() to plot only the MLE-recalibrated and 95% boldness-recalibrated sets as see in the right plot in Figure 2.
> lineplot(df=lp\$df, plot_original=FALSE, t_levels=c(0.95))
Figure 2: Demonstration of the lineplot() function visualizing how predictions change from varying recalibration strategies. On the left plot, the first column of points is the original set of probability predictions. The second column are the MLE-recalibrated predictions (i.e., after recalibrating with $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$). The last three columns are the predictions after 95%, 90%, and 80% boldness-recalibration (i.e. $\mathbf{x}_{0.95}$, $\mathbf{x}_{0.90}$, and $\mathbf{x}_{0.80}$) respectively. On the right plot, the columns are the MLE-recalibrated and 95% boldness-recalibrated probabilities. A line is used to connect each original prediction to where it ends up after each recalibration procedure. Blue points and lines correspond to predictions for which there was a foreclosure (i.e. “events”). Red corresponds to transactions in which there was no foreclosure (i.e. “non-events”). Achieved $P(M_{c}|\mathbf{y})$s are reported on the x-axis in parentheses.
There are a few key takeaways from the visualizations in Figure 2. First, the original foreclosure predictions are relatively accurate in terms of classification accuracy as more of the observations are colored blue (corresponding to foreclosure) for higher probabilities and more are colored red (corresponding to non-foreclosure) for lower probabilities. In fact, these predictions achieve an area under the ROC curve (AUC) of 0.8. Second, the original predictions are substantially biased, as they consistently underestimating the rate of foreclosure and MLE-recalibration corrects this bias. This is mostly likely due to using a model trained on pre-recession data from 2005 to estimate foreclosure rates for transactions from post-recession in 2010. Finally, these predictions can also benefit from being spread out under boldness-recalibration, as their spread has increased from the original set to the 95% boldness-recalibrated set. This is the maximum boldness achievable while maintaining a calibration probability of 0.95, subject to classification accuracy.
Next, we will use the plot_params() function to see the posterior model probability surface. As described in 3.6, we need to specify bounds for $\delta$ and $\gamma$ over which we want to plot and the density of the grid of values. Arguments dlim and glim take the bounds for $\delta$ and $\gamma$ over which the user wishes to plot. Argument k is used to specify a value for k. In the code below, we examine k=200 values of $\delta\in[8.5,15]$ and 200 values of $\gamma\in[1.05,1.5]$. Since we are interested in 95, 90, and 80% boldness-recalibration, we will also use t_levels=c(0.95, 0.9, 0.8) to specify that we want contours drawn at those three levels. Additionally, since this plot uses base R graphics, we can add points denoting the boldness-recalibration parameters and MLEs using points(). Recall, these parameter estimates are saved in br95, br90, br80, and bt. The result of the code below to generate the lineplot is shown in Figure 3
> plot_params(foreclosure\$x, foreclosure\$y,
+              dlim = c(8.5, 15), glim = c(1.05, 1.5), len.out=200,
+              t_levels = c(0.95, 0.9, 0.8))
> # add 95% B-R params
> points(br95\$BR_params[1], br95\$BR_params[2], pch=19, col="white")
> # add 90% B-R params
> points(br90\$BR_params[1], br90\$BR_params[2], pch=19, col="white")
> # add 80% B-R params
> points(br80\$BR_params[1], br80\$BR_params[2], pch=19, col="white")
> # add MLEs
> points(bt\$MLEs[1], bt\$MLEs[2], pch=4, col="white", cex=2)
Figure 3: Color contour plot demonstrating the plot_params() function. The points in white along the contours denote the boldness-recalibration parameters. The $\times$ denotes the MLEs.
The key takeaways from this visualization are the locations of high posterior model probability and the contours denoting the regions of interest for boldness-recalibration. This plot can aid users is validating the MLEs and boldness-recalibration estimates for $\delta$ and $\gamma$. For example, if the optimizer provides boldness-recalibration estimates that lie outside of the corresponding contour on the plot, users may need to make changes to the optimization routine to achieve convergence. Similarly, users should be concerned if the MLEs do no fall roughly in the middle of the dark red (high probability of calibration) region. In this cast study, the boldness-recalibration parameters are located along their corresponding contour. While the optimization is not restricted to searching only directly on the contour of $P(M_{c}|\mathbf{y})=t$, the optimal boldness-recalibration parameters all fall on their corresponding contours. A reason for this is that $\gamma$ is the scale parameter on the log-odds scale. Thus values of $\gamma$ that are further from 0 produce more bold predictions. Additionally, this is a direct result of the tension between calibration and boldness. As we increase $\gamma$ beyond $\hat{\gamma_{MLE}}$, i.e. more away from maximal calibration marked by the $\times$ (the location of the MLEs), we gain boldness. Naturally, this places the solution to our optimization at the boundary of the constraint which is visualized by the contours.
5 Conclusion
The BRcal package offers a variety of tools for assessing and improving probability predictions. One key contribution of this R package is the brcal() function which allows users to maximize the boldness (spread) in predictions while maintaining a specified level of calibration as measured by the posterior model probability of calibration. The other main contributions of the BRcal package are (i) MLE-recalibration, (iii) calibration assessment, and (iv) supporting visualizations. The BRcal package, with documentation and a user-friendly tutorial are available to users on CRAN and Github. Overall, the BRcal package is a useful tool to achieve more actionable predictions better decision making.
Conflict of Interest On behalf of all authors, the corresponding author states that there is no conflict of interest.
Appendix A Derivation of Gradients
This appendix shows the derivation of the gradients shown in Equations (13) and (15) in Sections A.2 and A.3 using related intermediate results presented in Section A.1.
A.1 Intermediate Results
The intermediate results in this section are useful for more clearly and concisely writing the derivations to follow. Let $x_{i}^{\prime}=c(x_{i};\delta,\gamma)=\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$ and $\hat{\delta}=\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}=\hat{\gamma}_{MLE}^{\prime}$ be the MLEs for $x_{i}^{\prime}$. We redefine these values for conciseness in derivations throughout the appendices. Note $x_{i,MLE}^{\prime}=c(x_{i}^{\prime};\hat{\delta},\hat{\gamma})$, which is equivalent to $c(x_{i};\hat{\delta}\delta^{\hat{\gamma}},\gamma\hat{\gamma})$ by A.1.1.
A.1.1 Nested LLO-Adjustment
Consider any LLO-adjusted prediction $x_{i}$ by $\delta$ and $\gamma$ that is then LLO-adjusted again by different parameters $\delta^{\prime}$ and $\gamma^{\prime}$. The resulting prediction can be re-written as follows in terms of a single LLO-adjustment using both sets of parameters.
$c(c(x_{i};\delta,\gamma);\delta^{\prime},\gamma^{\prime})$ $=\frac{\delta^{\prime}(c(x_{i};\delta,\gamma))^{\gamma^{\prime}}}{\delta^{\prime}(c(x_{i};\delta,\gamma))^{\gamma^{\prime}}+(1-c(x_{i};\delta,\gamma))^{\gamma^{\prime}}}$
$=\frac{\delta^{\prime}\left(\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}\right)^{\gamma^{\prime}}}{\delta^{\prime}\left(\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}\right)^{\gamma^{\prime}}+\left(1-\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}\right)^{\gamma^{\prime}}}$
$=\frac{\frac{\delta^{\prime}\delta^{\gamma^{\prime}}x_{i}^{\gamma\gamma^{\prime}}}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})^{\gamma^{\prime}}}}{\frac{\delta^{\prime}\delta^{\gamma^{\prime}}x_{i}^{\gamma\gamma^{\prime}}}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})^{\gamma^{\prime}}}+\frac{(1-x_{i})^{\gamma\gamma^{\prime}}}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})^{\gamma^{\prime}}}}$
$=\frac{\delta^{\prime}\delta^{\gamma^{\prime}}x_{i}^{\gamma\gamma^{\prime}}}{\delta^{\prime}\delta^{\gamma^{\prime}}x_{i}^{\gamma\gamma^{\prime}}+(1-x_{i})^{\gamma\gamma^{\prime}}}$
$=c(x_{i};\delta^{\prime}\delta^{\gamma^{\prime}},\gamma,\gamma^{\prime})$
A.1.2 Relationship Between LLO-adjusted $x_{i}$ and $1-x_{i}$
As mentioned in [1], if LLO-adjusted predicted probability for an “event” is defined by $x_{i}^{\prime}=c(x_{i};\delta,\gamma)$, then the corresponding LLO-adjusted prediction for a “non-event” is defined by $1-x_{i}^{\prime}=1-c(x_{i};\delta,\gamma)$. This can be re-written as the following LLO-adjustment.
$1-x_{i}^{\prime}$ $=1-c(x_{i};\delta,\gamma)$
$=1-\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=\frac{(1-x_{i})^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=\frac{\frac{1}{\delta}(1-x_{i})^{\gamma}}{\frac{1}{\delta}(1-x_{i})^{\gamma}+x_{i}^{\gamma}}$
$=c(1-x_{i};\frac{1}{\delta},\gamma)$
A.1.3 Log LLO-Adjustment
The following results pertains to log LLO-adjusted predictions needed for calculations of the gradients in Sections A.2 and A.3.
$log(x_{i}^{\prime})$ $=log(\delta)+\gamma log(x_{i})-log(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})$
$log(1-x_{i}^{\prime})$ $=log\left(c(1-x_{i};\frac{1}{\delta},\gamma)\right)$
$=\gamma log(1-x_{i})-log\left((1-x_{i})^{\gamma}+\delta x_{i}^{\gamma}\right)$
$log(x_{i,MLE}^{\prime})$ $=log(\hat{\delta}\delta^{\hat{\gamma}})+\gamma\hat{\gamma}log(x_{i})-log(\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}})$
$=log(\hat{\delta})+{\hat{\gamma}}log(\delta)+\gamma\hat{\gamma}log(x_{i})-log(\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}})$
$log(1-x_{i,MLE}^{\prime})$ $=\gamma\hat{\gamma}log(1-x_{i})-log\left((1-x_{i})^{\gamma\hat{\gamma}}+\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}\right)$
A.1.4 Intermediate Partial Derivatives
The following results pertain to derivatives needed for calculations of the gradients in Sections A.2 and A.3.
$\frac{\partial}{\partial\delta}x_{i}^{\prime}$ $=\frac{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})x_{i}^{\gamma}-\delta x_{i}^{2\gamma}}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})^{2}}$
$=\frac{x_{i}^{\gamma}(1-x_{i})^{\gamma}}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})}$
$=\frac{1}{\delta}\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}\frac{(1-x_{i})^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=\frac{1}{\delta}c(x_{i};\delta,\gamma)c(1-x_{i};\frac{1}{\delta},\gamma)$
$=\frac{1}{\delta}x_{i}^{\prime}(1-x_{i}^{\prime})$
$\frac{\partial}{\partial\gamma}x_{i}^{\prime}$ $=\frac{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})\delta x_{i}^{\gamma}log(x_{i})-\delta x_{i}^{\gamma}(\delta x_{i}^{\gamma}log(x_{i})+(1-x_{i})^{\gamma}log(1-x_{i}))}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})^{2}}$
$=\frac{\delta x_{i}^{\gamma}(1-x_{i})^{\gamma}}{(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma})}log\left(\frac{x_{i}}{1-x_{i}}\right)$
$=c(x_{i};\delta,\gamma)c(1-x_{i};\frac{1}{\delta},\gamma)log\left(\frac{x_{i}}{1-x_{i}}\right)$
$=x_{i}^{\prime}(1-x_{i}^{\prime})log\left(\frac{x_{i}}{1-x_{i}}\right)$
$\frac{\partial}{\partial\delta}log(x_{i}^{\prime})$ $=\frac{1}{\delta}-\frac{x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=\frac{(1-x_{i})^{\gamma}}{\delta\left(\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}\right)}$
$=\frac{1}{\delta}c(1-x_{i};\frac{1}{\delta},\gamma)$
$=\frac{1}{\delta}(1-x_{i}^{\prime})$
$\frac{\partial}{\partial\gamma}log(x_{i}^{\prime})$ $=log(x_{i})-\frac{\delta x_{i}^{\gamma}log(x_{i})+(1-x_{i})^{\gamma}log(1-x_{i})}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=\frac{(1-x_{i})^{\gamma}log\left(\frac{x_{i}}{1-x_{i}}\right)}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=log\left(\frac{x_{i}}{1-x_{i}}\right)c(1-x_{i};\frac{1}{\delta},\gamma)$
$=log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i}^{\prime})$
$\frac{\partial}{\partial\delta}log(1-x_{i}^{\prime})$ $=-\frac{x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=-\frac{1}{\delta}\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=-\frac{1}{\delta}x_{i}^{\prime}$
$\frac{\partial}{\partial\gamma}log(1-x_{i}^{\prime})$ $=log(1-x_{i})-\frac{log(1-x_{i})(1-x_{i})^{\gamma}+log(x_{i})\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=log\left(\frac{1-x_{i}}{x_{i}}\right)\frac{\delta x_{i}^{\gamma}}{\delta x_{i}^{\gamma}+(1-x_{i})^{\gamma}}$
$=log\left(\frac{1-x_{i}}{x_{i}}\right)x_{i}^{\prime}$
$=-log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime}$
$\frac{\partial}{\partial\delta}log(x_{i,MLE}^{\prime})$ $=\frac{\hat{\gamma}}{\delta}-\frac{\hat{\delta}\hat{\gamma}\delta^{\hat{\gamma}-1}x_{i}^{\gamma\hat{\gamma}}}{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}}}$
$=\frac{\hat{\gamma}(1-x_{i})^{\gamma\hat{\gamma}}}{\delta(\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}})}$
$=\frac{\hat{\gamma}}{\delta}c(1-x_{i};\frac{1}{\hat{\delta}\delta^{\hat{\gamma}}},\gamma\hat{\gamma})$
$=\frac{\hat{\gamma}}{\delta}(1-x_{i,MLE}^{\prime})$
$\frac{\partial}{\partial\gamma}log(x_{i,MLE}^{\prime})$ $=\hat{\gamma}log(x_{i})-\frac{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}log(x_{i})\hat{\gamma}+(1-x_{i})^{\gamma\hat{\gamma}}log(1-x_{i})\hat{\gamma}}{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}}}$
$=\frac{\hat{\gamma}log(\frac{x_{i}}{1-x_{i}})(1-x_{i})^{\gamma\hat{\gamma}}}{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}}}$
$=\hat{\gamma}log\left(\frac{x_{i}}{1-x_{i}}\right)c(1-x_{i};\frac{1}{\hat{\delta}\delta^{\hat{\gamma}}},\gamma\hat{\gamma})$
$=\hat{\gamma}log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i,MLE}^{\prime})$
$\frac{\partial}{\partial\delta}log(1-x_{i,MLE}^{\prime})$ $=-\frac{\hat{\delta}\hat{\gamma}\delta^{\hat{\gamma}-1}x_{i}^{\gamma\hat{\gamma}}}{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}}}$
$=-\frac{\hat{\gamma}}{\delta}\frac{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}}{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}}}$
$=-\frac{\hat{\gamma}}{\delta}x_{i,MLE}^{\prime}$
$\frac{\partial}{\partial\gamma}log(1-x_{i,MLE}^{\prime})$ $=\hat{\gamma}log(1-x_{i})-\frac{(1-x_{i})^{\gamma\hat{\gamma}}log(1-x_{i})\hat{\gamma}+\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}log(x_{i})\hat{\gamma}}{(1-x_{i})^{\gamma\hat{\gamma}}+\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}}$
$=\frac{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}\hat{\gamma}log(\frac{1-x_{i}}{x_{i}})}{\hat{\delta}\delta^{\hat{\gamma}}x_{i}^{\gamma\hat{\gamma}}+(1-x_{i})^{\gamma\hat{\gamma}}}$
$=\hat{\gamma}log(\frac{1-x_{i}}{x_{i}})c(x_{i};\hat{\delta}\delta^{\hat{\gamma}},\gamma\hat{\gamma})$
$=\hat{\gamma}log(\frac{1-x_{i}}{x_{i}})x_{i,MLE}^{\prime}$
$=-\hat{\gamma}log(\frac{x_{i}}{1-x_{i}})x_{i,MLE}^{\prime}$
A.2 Objective Function Gradients
The following series of equations show the derivation of the gradients of the objective function in Equation 10 using the results from Section A.1.
$\frac{\partial f(\delta,\gamma)}{\partial\delta}$ $=-\frac{\frac{\partial}{\partial\delta}\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})^{2}}{2\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})^{2}}}$
$=-\frac{\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\frac{\partial}{\partial\delta}(x_{i}^{\prime}-\bar{x}^{\prime})}{(n-1)sd(\mathbf{x^{\prime}})}$
$=-\frac{\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(\frac{1}{\delta}x_{i}^{\prime}(1-x_{i}^{\prime})-\frac{1}{n}\sum_{j=1}^{n}\frac{1}{\delta}x_{j}^{\prime}(1-x_{j}^{\prime})\right)}{(n-1)sd(\mathbf{x^{\prime}})}$
$=-\frac{\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(x_{i}^{\prime}(1-x_{i}^{\prime})-\frac{1}{n}\sum_{j=1}^{n}x_{j}^{\prime}(1-x_{j}^{\prime})\right)}{\delta(n-1)sd(\mathbf{x^{\prime}})}$
$\frac{\partial f(\delta,\gamma)}{\partial\gamma}$ $=-\frac{\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\frac{\partial}{\partial\gamma}(x_{i}^{\prime}-\bar{x}^{\prime})}{(n-1)sd(\mathbf{x^{\prime}})}$
$=-\frac{\sum_{i=1}^{n}(x_{i}^{\prime}-\bar{x}^{\prime})\left(x_{i}^{\prime}(1-x_{i}^{\prime})log\left(\frac{x_{i}}{1-x_{i}}\right)-\frac{1}{n}\sum_{j=1}^{n}x_{j}^{\prime}(1-x_{j}^{\prime})log\left(\frac{x_{j}}{1-x_{j}}\right)\right)}{(n-1)sd(\mathbf{x^{\prime}})}$
A.3 Constraint Function Gradients
The following series of equations show the derivation of the gradients of the constraint function in Equation 12 using the results from Section A.1.
$\frac{\partial f(\delta,\gamma)}{\partial\delta}$ $=\frac{1}{\left(1+BF\frac{P(M_{u})}{P(M_{c})}\right)^{2}}\frac{\partial}{\partial\delta}\left(1+BF\frac{P(M_{u})}{P(M_{c})}\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})^{2}\frac{P(M_{u})}{P(M_{c})}exp\left\{-\frac{1}{2}(BIC_{u}-BIC_{c})\right\}\frac{\partial}{\partial\delta}-\frac{1}{2}(BIC_{u}-BIC_{c})$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})^{2}\left(\frac{1}{P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})}-1\right)$
$\frac{\partial}{\partial\delta}\left(-log(n)+\sum_{i=1}^{n}y_{i}log(x_{i,MLE}^{\prime})+(1-y_{i})log(1-x_{i,MLE}^{\prime})-y_{i}log(x_{i}^{\prime})-(1-y_{i})log(1-x_{i}^{\prime})\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))$
$\left(\sum_{i=1}^{n}y_{i}\frac{\hat{\gamma}}{\delta}(1-x_{x_{i},MLE}^{\prime})+(1-y_{i})(-\frac{\hat{\gamma}}{\delta}x_{i,MLE}^{\prime})-y_{i}\frac{1}{\delta}(1-x_{i}^{\prime})-(1-y_{i})(-\frac{1}{\delta}x_{i}^{\prime})\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\left(\sum_{i=1}^{n}\frac{\hat{\gamma}}{\delta}y_{i}-\frac{\hat{\gamma}}{\delta}x_{i,MLE}^{\prime}-\frac{1}{\delta}y_{i}+\frac{1}{\delta}x_{i}^{\prime}\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\frac{1}{\delta}\left(\sum_{i=1}^{n}(\hat{\gamma}-1)y_{i}-\hat{\gamma}x_{i,MLE}^{\prime}+x_{i}^{\prime}\right)$
$\frac{\partial f(\delta,\gamma)}{\partial\gamma}$ $=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))$
$\frac{\partial}{\partial\gamma}\left(-log(n)+\sum_{i=1}^{n}y_{i}log(x_{i,MLE}^{\prime})+(1-y_{i})log(1-x_{i,MLE}^{\prime})-y_{i}log(x_{i}^{\prime})-(1-y_{i})log(1-x_{i}^{\prime})\right)$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}\hat{\gamma}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i,MLE}^{\prime})$
$-\hat{\gamma}(1-y_{i})log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i,MLE}^{\prime}-y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)(1-x_{i}^{\prime})+(1-y_{i})log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime}\bigg{)}$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}\hat{\gamma}y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)-\hat{\gamma}log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i,MLE}^{\prime}$
$-y_{i}log\left(\frac{x_{i}}{1-x_{i}}\right)+log\left(\frac{x_{i}}{1-x_{i}}\right)x_{i}^{\prime}\bigg{)}$
$=P(M_{c}|\mathbf{y},\mathbf{x}^{\prime})(1-P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}))\bigg{(}\sum_{i=1}^{n}log\left(\frac{x_{i}}{1-x_{i}}\right)\big{(}(\hat{\gamma}-1)y_{i}-\hat{\gamma}x_{i,MLE}^{\prime}+x_{i}^{\prime}\big{)}\bigg{)}$
Appendix B Strategy for Specifying Parameter Grid in plot_params()
When using plot_params(), some care is required for the specifications of arguments dlim, glim, and k, which together determine the fineness of your grid of parameter values plotted. We suggest the following scheme for determining the grid of parameter values to search. First, fix k at some small number, less than 20 for sake of computation time. Then, center a grid with small range around $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$ for $\mathbf{x}$. Increase the size of k until your grid detects approximated the probability of calibration at the MLEs that you expect. At this point, you may find that your grid does not fully encompass the range of parameters values which produce recalibrated probabilities with meaningfully high posterior probability of being calibrated. If this is the case, expand your grid until it the region with high probability of calibration is covered. Then, increase k to create a fine grid of values. Additionally, we caution users from including $\gamma=0$ in the grid. This setting recalibrates all values in $\mathbf{x}$ to a single value which is not desirable in practice. Unless the singular value is near the base rate, the set will be poorly calibrated and minimally bold, which does not align with the goal of boldness-recalibration.
As mentioned, the computation time to construct this plot can be high if the grid of parameter values is very fine or the sample size of predictions is very large. In an effort to reduce this bottleneck, we have developed as strategy to circumvent unnecessary calls to the optim(). To calculate $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$ for each grid cell, this requires evaluating the likelihood at the MLEs for each LLO-adjusted set $\mathbf{x}^{\prime}_{ij}$, which in turn requires a call to optim() to get the MLEs. However, given the MLEs for the original set $\mathbf{x}$ and the $\delta_{i}$ and $\gamma_{j}$ from the grid, we show that it is possible to get the MLEs for $\mathbf{x}^{\prime}_{ij}$ without optimization via the result in Appendix C. Thus, we can reduce the number of optim() calls from one call for each grid cell to just one call overall to find the MLEs for the original set. An example of a direct result of this is reducing the computation time over a grid for n=868, k=100 from approximately 2 hours to under a minute.
Appendix C plot_params() Speed Up
In the plot produced by plot_params(), recall each grid cell is color coded by
$P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$ $=\left(1+\frac{P(M_{u})}{P(M_{c})}exp\{-\frac{1}{2}(BIC_{u}-BIC_{c})\}\right)^{-1}$
where $BIC_{u}=2\times log(n)-2\times log(\pi(\hat{\delta}_{MLE},\hat{\gamma}_{MLE}|\mathbf{x}_{ij}^{\prime},\mathbf{y})))$, $\mathbf{x}_{ij}^{\prime}=c(\mathbf{x};\delta_{i},\gamma_{j})$, and $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ are the MLEs for $\mathbf{x}_{ij}^{\prime}$. Notice that $BIC_{u}$ requires a unique calculation of $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ for each $i,j$. In this section, we show that $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ can be calculated using only the MLEs for $\mathbf{x}$ ($\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$), and the grid shift parameters ($\delta_{i}$ and $\gamma_{j}$). This is important as it circumvents an additional call to optim() for each grid cell.
First, note that
$(\hat{\delta}_{MLE},\hat{\gamma}_{MLE})=argmax_{(\hat{\delta},\hat{\gamma})}\left(\prod_{k=1}^{n}c(x_{k};\hat{\delta},\hat{\gamma})^{y_{k}}(1-c(x_{k};\hat{\delta},\hat{\gamma}))^{1-y_{k}}\right).$ (17)
Similarly,
$(\hat{\delta}_{MLE}^{\prime},\hat{\gamma}_{MLE}^{\prime})$ $=argmax_{(\hat{\delta}^{\prime},\hat{\gamma}^{\prime})}\left(\prod_{k=1}^{n}c(x_{k}^{\prime};\hat{\delta}^{\prime},\hat{\gamma}^{\prime})^{y_{k}}(1-c(x_{k}^{\prime};\hat{\delta}^{\prime},\hat{\gamma}^{\prime}))^{1-y_{k}}\right)$ (18)
$=argmax_{(\hat{\delta}^{\prime},\hat{\gamma}^{\prime})}\left(\prod_{k=1}^{n}c(c(x_{k};\delta_{i},\gamma_{j});\hat{\delta}^{\prime},\hat{\gamma}^{\prime})^{y_{k}}(1-c(c(x_{k};\delta_{i},\gamma_{j});\hat{\delta}^{\prime},\hat{\gamma}^{\prime}))^{1-y_{k}}\right)$ (19)
$=argmax_{(\hat{\delta}^{\prime},\hat{\gamma}^{\prime})}\left(\prod_{k=1}^{n}c(x_{k};\hat{\delta}^{\prime}\delta_{i}^{\hat{\gamma}^{\prime}},\hat{\gamma}^{\prime}\gamma_{j})^{y_{k}}(1-c(x_{k};\hat{\delta}^{\prime}\delta_{i}^{\hat{\gamma}^{\prime}},\hat{\gamma}^{\prime}\gamma_{j}))^{1-y_{k}}\right)$ (20)
by the result in Appendix A.1.1. This implies that when both Equation (17) and (18) are maximized, $\hat{\delta}_{MLE}=\hat{\delta}_{MLE}^{\prime}\delta_{i}^{\hat{\gamma}_{MLE}^{\prime}}$ and $\hat{\gamma}_{MLE}=\hat{\gamma}_{MLE}^{\prime}\gamma_{j}$. Solving for $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ results in
$\hat{\delta}_{MLE}^{\prime}$ $=\frac{\hat{\delta}_{MLE}}{\delta_{i}^{\hat{\gamma}_{MLE}/\gamma_{j}}}$ (21)
$\hat{\gamma}_{MLE}^{\prime}$ $=\frac{\hat{\gamma}_{MLE}}{\gamma_{j}}.$ (22)
Using these expression, we circumvent the need to use optim() to optimize Equation (18) to get $\hat{\delta}_{MLE}^{\prime}$ and $\hat{\gamma}_{MLE}^{\prime}$ for each grid cell. Note that the expressions for $\hat{\gamma}_{MLE}^{\prime}$ is undefined when $\gamma_{j}=0$, which is a valid grid cell value. In this case, all probability are set to 0.5, regardless of $\delta_{i}$.